OpenStudy (anonymous):

Find the radius of convergence of the series:

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ x^n }{ n9^n }\]

OpenStudy (anonymous):

Already did ratio test: \[\lim_{n \rightarrow \infty} =\frac{ |x| }{ 9 }\], therefore the series converges when |x|/9=0 I already have my endpoints done and checked, but I'm having trouble finding R

OpenStudy (amistre64):

x^(n+1) (n) 9^(n) ------------------- x^(n) (n+1) 9^(n+1) x (n) ------ 9 (n+1) |x/9| lim n/(n+1) = |x/9| the radius of convergence is half the interval of convergence |x/9| < 1 -1 < x/9 < 1 -9 < x < 9 since the interval is 18, whats the radius?

OpenStudy (amistre64):

and to check the endpoints, just let x=-9, or 9 and see if it limits